# HARMONIOUS – Human-like reactive motion control and multimodal perception for humanoid robots

### v1

@@ II-F Obstacle processing
The obstacle virtually exists for a specific amount of time with a decreasing threat level imitating that the obstacle is moving away (similar to [9]) to prevent dangerous and jerky movements and to give humans time to react when sensor data are missing or for a post-collision reaction. For example, the threat disappears once the robot moves away from the tactile collision. Without virtual preservation of the obstacle, the robot would return to its original position even when the obstacle is still there.
Visual obstacles in our experiments are human keypoints detected by an RGB-D camera. The method for their detection is based on the work by Docekal et al. [34]. The detected obstacles (see a schematic illustration in Fig. 2(a)) are mapped to parts of the skin where physical contact would eventually occur using PPS projection, similar to [30]. All these mappings are combined into a single position per body part. The collision direction is the vector from the clustered position to the obstacle and the threat level is calculated from their distance.
Proximity sensor data (see a schematic illustration in Fig. 2(b)) are not projected onto the skin taxel, as these sensors are placed to cover parts of the robot without the skin; thus, there is no skin taxel to which information can be projected. However, they are projected onto the sensor position. The threat level is computed from the measured distance, and the collision direction is the direction vector of the sensor beam.
Tactile obstacles (see a schematic illustration in Fig. 2(c)) are processed in the same way as in [17]. Contacts detected by artificial skin sensors (taxels) with a pressure above the threshold value are clustered to super contacts to diminish spiking effects and eliminate false positive detections. The threat level corresponds to the highest measured pressure in the neighborhood, and the collision direction is the normal vector of the central taxel.
II-G Collision avoidance parameters
Obstacle avoidance constraints (Eq. 12) are parameterized by two positive parameters—additive $k_{1}$ and multiplicative $k_{2}$. The parameter $k_{1}$ defines when the maximum approach velocity becomes the minimum repulsive velocity, that is, when $\mathbf{\dot{x}_{a}}$ is lower than zero. The parameter $k_{1}$ has a constant value of 0.3 in our experiments. The parameter $k_{2}$ scales the maximum approach velocity. We found that body links cannot move with the same minimum repulsive velocity due to the kinematic constitution of the robot. The same scale for all parts leads to unfeasible constraints for the upper arm and torso (high value) or small and slow avoidance for the hand and forearm (low value). From our observation, $k_{2}$ should increase from the torso to the hand. In our experiments, $k_{2}$ is 0.06, 0.06, 0.33, 0.53 for the body parts of the torso, upper arm, forearm, and hand, respectively.
II-H Self-collision avoidance
@@
We find the closest pair of virtual obstacles and projected positions for each combination of body parts. If their distance $d_{C}$ is less than a threshold, we add the virtual obstacle to the set of obstacles with a threat level $a_{t}=1-50d_{C}/3$. The collision point is the projected position of the obstacle on the robot ($\mathbf{P}_{C}$), the collision direction ($\mathbf{n}_{C}$) is the vector from the collision point to the virtual obstacle, and the gain ($V_{C}$) is equal to 1.2. Finally, we compute the linear inequality constraint for these obstacles in the same way as for the external obstacles.
II-I Static obstacles
@@
Sometimes, the experiment contains known static objects that the robot should avoid. The almost same procedure as for self-collision avoidance can also be used in this case. Virtual obstacles are sampled from the static object instead of the robot body. The projected positions are points in the robot body parts that can collide with the object. This approach was used, for example, to avoid the table during the interactive game demonstration.
II-J Relative position constraint
@@
In addition to the usual limitations, such as keeping a target pose, we can incorporate another task restriction into the problem as (in-)equality constraints. For the bimanual task experiment, we added one constraint per each axis to keep a specified relative position between the hands (end effectors). The constraint can be written in the vector form as
$\mathbf{x^{p}}-\mathbf{x^{s}}=\mathbf{d_{\mathrm{rel}}}$
@@
where $\mathbf{x^{p/s}}$ are the positions of the end effectors, and $\mathbf{d_{\mathrm{rel}}}$ is a relative position vector. This equation can be reformulated for our QP problem with joint velocities as
$\begin{split}\mathbf{x_{0}^{p}}+\begin{bmatrix}\mathbf{J_{t,pos}^{p}}(\mathbf{q^{t}})&\mathbf{J_{a,pos}^{p}}(\mathbf{q^{p}})\end{bmatrix}\begin{bmatrix}\mathbf{\dot{q}^{t}}\\ \mathbf{\dot{q}^{p}}\end{bmatrix}&-\\ \mathbf{x_{0}^{s}}-\begin{bmatrix}\mathbf{J_{t,pos}^{s}}(\mathbf{q^{t}})&\mathbf{J_{a,pos}^{s}}(\mathbf{q^{s}})\end{bmatrix}\begin{bmatrix}\mathbf{\dot{q}^{t}}\\ \mathbf{\dot{q}^{s}}\end{bmatrix}&=\mathbf{d_{\mathrm{rel}}}\end{split}$ (14)
@@
where $\mathbf{x_{0}^{p/s}}$ are the current positions of the end effectors, $\mathbf{J^{p/s}_{a,pos}}(\mathbf{q^{p/s}}),\mathbf{J^{p/s}_{t,pos}}(\mathbf{q^{t}})$ are the translational velocity components of the Jacobian matrices for arm and torso, respectively.
II-K Joint position constraints
@@
The joint position bounds for the QP problem have to be set in such a way that the robot joint limits are not exceeded. In our solution, we are inspired by the shaping policy for avoiding joint limits in [4] to maximize reachability in the arm workspace. It consists of a flat region in most of the joint range. Nevertheless, we replaced the original hyperbolic tangent functions by linear and constant functions to fulfill the constraint condition of the QP problem. Since the problem is solved at the velocity level, the joint position limits are expressed using the joint velocities for the joint $i$ as:
@@
III Experiments and Results
We prepared seven experiments with a simulated iCub robot and three with a real one to validate the performance of the proposed solution and compare it with other solutions. The experiments overview can be seen in Tab. II.
@@ 13 Real Bimanual
III-A Experiments in simulation
In simulation, we evaluated the performance of the controller (see Sec. II-B) and the trajectory sampling (see Sec. II-D). The first set of experiments aims to assess reachability, movement smoothness, and reference tracking, without obstacles nearby. Reachability is tested in a reaching experiment (Exp 1) with the targets taken from a grid of 3x3x3 positions in Cartesian space in random order and randomly switched two orientations. The positions are the Cartesian product of $x=$ [-0.23, -0.19, -0.15] m, $y=$ [0.11, 0.15, 0.19] m, $z=$ [0.08, 0.12, 0.16] m, and the orientations are $o_{1}=$ [-0.15, -0.79, 0.59, 3.06], $o_{2}=$ [-0.11, 0.99, 0.02, 3.14] in angle-axis notation. Smooth movement is evaluated in a point-to-point experiment (Exp 2), where the robot moves between two alternating poses $p_{1}=$ [-0.23, 0.26, 0.02, -0.15, -0.79, 0.59, 3.06] and $p_{2}=$ [-0.26, 0.03, 0.03, -0.11, 0.99, 0.01, 3.14], where the first three values are position coordinates in meters and the last four values are the orientation in angle-axis notation. Finally, a circular movement experiment (Exp 3) is chosen to validate the reference tracking. The overview of the experiments can be found in Tab. II.
In addition to the proposed solution, other alternatives and state-of-the-art methods are also evaluated. Specifically, we compare — 1) HARMONIOUS – the proposed solution in the single-arm version; 2) HARMONIOUS2 – the proposed solution in the dual-arm version when the goal of the other arm is to hold a specified position; 3) NEO – the controller by Haviland and Corke [8]; 4) React by Ngyuen et al. [17]; 5) Cart – Cartesian controller for the iCub robot [4], which is evaluated only in Exp 1 and Exp 2 as it primarily serves for point-to-point movements and does not feature avoidance constraints. Controllers 1–4 are evaluated with and without local trajectory sampling (LTS), described in Sec. II-D. Cart has its own local trajectory optimization.
@@ III-A Experiments in simulation
TABLE III: Reachability experiment (Exp 1). Success rate of combined reachability in position and orientation. LTS stands for local trajectory sampling. Cart has its own local trajectory optimization.
The results for the reachability experiment (Exp 1) are shown in Tab. III. In this experiment, our controller has the highest number of targets that are reached in 6D pose (i.e., position and orientation). The target is reached when the end effector is, at the same time, closer than 5 mm and 0.1 rad to the target. The impact of the local trajectory sampling is visible as all controllers have better (or at least the same) reachability results with sampling than without sampling.
The results of the movement smoothness experiment (Exp 2) show the main impact of the local trajectory sampling (see Fig. 3(a)). Position sampling (Fig. 3(a) left) creates a bell-shaped Cartesian translation velocity profile typical for a human-arm motion [50]. The velocity profile of Cart is the most similar to the exact minimum-jerk movement, as it is specially optimized for it. Other versions are almost identical; only React does not reach zero velocity after 5 s. All versions without local sampling have snap onset and slow jerky decay. Orientation sampling (Fig. 3(a) right) creates a torque-minimal path for the rotation as it travels along the straightest path of the rounded surface of a sphere and it keeps the rotational velocity constant during movement [52], resulting in a visually smooth transition between two rotations. HARMONIOUS and NEO have a smooth angular velocity profile; React velocity profile is not smooth and does not reach zero after 5 s. Cart uses the same sampling for orientation as for position, and thus the resulting angular velocity profile is similar to that for the translational velocity.
@@
The next experiment is a circular movement experiment with both arms in opposite directions (Exp 4). It shows how the dual-arm solution works, particularly arm prioritization and self-collision avoidance. As the reference controllers are for single arms only (or separately), only our solution is shown (without trajectory sampling). Figure 3(c) shows the trajectories of both end effectors making circular movements. In this case, the right arm was selected as the “primary arm”. Therefore, the right arm successfully tracks the circle all the time, and the left arm prevents self-collision by keeping a safe distance instead of tracking the circle when the arms are too close.
@@
The last set of simulation experiments tests collision avoidance. For simplicity, the obstacle is virtual and is perceived without noise, omitting the influence of sensor data processing on the results. All three experiments are based on Exp 2 and they differ only in the direction of the obstacle — 1) the obstacle moves along the movement, i.e., in the y-axis direction (Exp 5-1); 2) the obstacle moves towards the robot, i.e., in the x-axis direction (Exp 5-2); 3) the obstacle falls in the z-axis direction (Exp 5-3). In this case, Cart is not included in the comparison because it cannot handle obstacles. React is also omitted because it was outperformed even in experiments without obstacles. Furthermore, more restrictive repulsive vectors were used instead of linear inequality constraints that allow motions in tangential directions. Therefore, we compare only HARMONIOUS and NEO. Figure 3(d) shows the fraction of targets reached during the experiment, the fraction of QP problems successfully solved during the experiments, and the distance between the obstacle and the closest point of the robot, where we assume that a distance of less than 25 mm is too dangerous. There is no significant difference between controllers when the obstacle moves along the robot’s movement (Exp 5-1). On the other hand, the obstacle moving towards the robot (Exp 5-2) is problematic for NEO. Half of the time, the targets are not reached within a time limit and the controller cannot solve the QP problem in almost 40 % runs. Furthermore, the small distances between the robot and the obstacle sometimes lead to collisions. As opposed to NEO, HARMONIOUS has no problem with collisions or reaching in this experiment. Our solution also works adequately in Exp 5-3. NEO did not reach all targets within the time limit, but the distance to an obstacle is safe.
III-B Real robot experiments – collision avoidance
@@
The experiments with the real robot were prepared to evaluate the obstacle processing part of the solution, described in Sec. II-F. We show collision avoidance (visual and proximity modalities) and post-collision reaction (tactile modality) and their combination while the robot keeps position (Exp 11) or moves back and forth between two points as in Exp 2 (Exp 12). When visual obstacles are involved, the iCub gaze controller [53] is employed to point the robot gaze to an appropriate position in the scene, thus compensating for the robot torso joint movements.
@@
Figure 5 graphically illustrates the experiment using snapshots. Figure 6 shows the operation of the main components of HARMONIOUS on plots, focusing on the left arm of the robot (“primary arm” in these experiments). Note that obstacles at or near the robot hand (visual obstacles near the hand or forearm, proximity at the back of the hand, or touch at the hand or forearm) inevitably compromise the task—staying close to the target—whereas visual obstacles or touch at more proximal body parts give the robot the possibility to exploit its kinematic redundancy and fulfill the task while simultaneously avoiding contacts. Parts of Exp 11—robot tasked with keeping the end effector position, while complying with constraints from obstacles spawned by the presence of the human—are shown in Fig. 5(a). In Fig. 5(a) (top), the robot starts in a steady state without obstacles around characterized by a zero target distance (first two seconds, green line). Once a proximity obstacle is detected, the end effector is moving away until the proximity obstacle constraint disappears from the controller and the target position is reached again. Physical contact of the human with the robot hand (around 5 s) and the forearm (around 7 and 9 s) generate constraints and trigger avoidance. Tactile collision with upper arm (12-13 s) and torso (14-15 s) generate avoidance but do not significantly compromise the task, as the robot kinematic redundancy is exploited by the controller.
@@
In Fig. 5(a) (center), the first 10 seconds are very similar to the top one. Touch on the forearm causes deviation from the target while contact with the upper arm is evaded while simultaneously staying close to the target. Touch at the hand (5 s) and proximity signal at the hand (7 s) cause deviation from the target. At around 12 s, a proximity obstacle at the hand is evaded but then avoidance stops as the robot is blocked by contact with the forearm. Once the touch disappears, the proximity collision avoidance continues. Finally, when the proximity obstacle disappears, a tactile stimulus on the hand causes a fast movement towards the target position.
@@
III-C Real robot – bimanual task
We prepared a bimanual task experiment (Exp 13) in which the robot holds a sponge in its hands (see Fig. 6(a)). The target is prescribed for the “primary arm” (left) and an additional constraint is introduced to keep the sponge between the two hands. At the same time, dynamic obstacles spawn additional constraints and avoidance. For this experiment, tactile inputs from the robot hands (palms) are ignored (the robot holds the sponge). This experiment aims to verify that the dual-arm version works and that other task constraints, such as the relative distance vector between the end effectors in this case, can be easily incorporated into the problem definition (the relative position constraint described in Sec. II-J).

### v2

@@ III-F Obstacle processing
To prevent jerky motions induced by vanishing obstacle readings, we apply a linear decay formula to simulate an obstacle slowly moving away from the robot [12]. Every obstacle virtually exists for a specific amount of time (called “surviving time”). For example, the threat disappears once the robot moves away from the tactile collision. Without virtual preservation of the obstacle, the robot would return to its original position even when the obstacle is still there.
Visual obstacles in our experiments are human keypoints detected by an RGB-D camera. The method for their detection is based on the work by Docekal et al. [36]. The detected obstacles (see a schematic illustration in Fig. 3(a)) are mapped to parts of the skin where a collision is most likely to occur using a PPS projection, similar to [4]. For each body part, we compute one visual obstacle as a weighted average of the mappings to that body part. The collision direction is the vector from the estimated contact position on the skin to the obstacle and the threat level is calculated from their distance.
Proximity sensor data (see a schematic illustration in Fig. 3(b)) are not projected onto skin taxels, as these sensors are placed to cover parts of the robot without the skin; therefore, there is no skin taxel to which information can be projected. However, they are projected onto the sensor position. The threat level is computed from the measured distance, and the collision direction is the direction vector of the sensor beam.
Tactile obstacles (see a schematic illustration in Fig. 3(c)) are processed in the same way as in [3]. Contacts detected by artificial skin sensors (taxels) with a pressure above the threshold value are clustered to super contacts (one per each body part) to diminish spiking effects and eliminate false positive detections. The threat level corresponds to the highest measured pressure in the neighborhood, and the collision direction is the normal vector of the central taxel.
III-G Collision avoidance parameters
Obstacle avoidance constraints (Eq. 12) are parameterized by two positive parameters—additive $k_{1}$ and multiplicative $k_{2}$. The parameter $k_{1}$ defines when the maximum approach velocity becomes the minimum repulsive velocity, that is, when $\dot{x}_{a}$ is lower than zero. The parameter $k_{1}$ has a constant value of 0.3 in our experiments. The parameter $k_{2}$ scales the maximum approach velocity. We found that body links cannot move with the same minimum repulsive velocity due to the kinematic constitution of the robot. The same scale for all parts leads to unfeasible constraints for the upper arm and torso (high value) or small and slow avoidance for the hand and forearm (low value). From our observation, $k_{2}$ should increase from the torso to the hand. In our experiments, $k_{2}$ is 0.06, 0.06, 0.33, 0.53 for the body parts of the torso, upper arm, forearm, and hand, respectively.
III-H Self-collision avoidance
@@
We find the closest pair of virtual obstacles and projected positions for each combination of body parts. If their distance $d_{C}$ is less than a threshold, we add the virtual obstacle to the set of obstacles with $a_{t}\cdot V_{C}=1.2-20d_{C}\geq 0$. The coefficients were empirically determined in simulation and on the real robot to obtain appropriate self-collision avoidance reactions. The collision point is the projected position of the obstacle on the robot ($\mathbf{P}_{C}$), the collision direction ($\mathbf{n}_{C}$) is the vector from the collision point to the virtual obstacle. Finally, we compute the linear inequality constraint for these obstacles in the same way as for the external obstacles.
III-I Static obstacles
@@
Sometimes, the experiment contains known static objects that the robot should avoid. The almost same procedure as for self-collision avoidance can also be used in this case. Virtual obstacles are sampled from the static object instead of the robot body. The projected positions are points on the robot body parts that are in danger of colliding with the object. This approach was used, for example, to avoid the table during the interactive game demonstration.
III-J Relative position constraint
@@
Other task restrictions can be incorporated into the problem as (in-)equality constraints. For the bimanual task experiment, we added one constraint per each axis to keep a specified relative position between the hands (end effectors). The constraint can be written in the vector form as
$\mathbf{x^{p}_{k+1}}-\mathbf{x^{s}_{k+1}}=\mathbf{d_{\mathrm{rel}}}$
@@
where $\mathbf{x^{p/s}_{k+1}}\in\mathbb{R}^{3}$ are the next (i.e. in timestep $k+1$) positions of the end effectors, and $\mathbf{d_{\mathrm{rel}}}\in\mathbb{R}^{3}$ is a relative position vector. This equation can be reformulated for our QP problem with joint velocities as
$\begin{split}\mathbf{x_{k}^{p}}+\begin{bmatrix}\mathbf{J_{t,pos}^{p}}(\mathbf{q^{t}})&\mathbf{J_{a,pos}^{p}}(\mathbf{q^{p}})\end{bmatrix}\begin{bmatrix}\mathbf{\dot{q}^{t}}\\ \mathbf{\dot{q}^{p}}\end{bmatrix}&-\\ \mathbf{x_{k}^{s}}-\begin{bmatrix}\mathbf{J_{t,pos}^{s}}(\mathbf{q^{t}})&\mathbf{J_{a,pos}^{s}}(\mathbf{q^{s}})\end{bmatrix}\begin{bmatrix}\mathbf{\dot{q}^{t}}\\ \mathbf{\dot{q}^{s}}\end{bmatrix}&=\mathbf{d_{\mathrm{rel}}}\end{split}$ (14)
@@
where $\mathbf{x_{k}^{p/s}}\in\mathbb{R}^{3}$ are the current (i.e. in timestep $k$) positions of the end effectors, $\mathbf{J^{p/s}_{a,pos}}(\mathbf{q^{p/s}})\in\mathbb{R}^{3\times(n-3)},\mathbf{J^{p/s}_{t,pos}}(\mathbf{q^{t}})\in\mathbb{R}^{3\times 3}$ are the translational velocity components of the Jacobian matrices for arm and torso, respectively.
III-K Joint position constraints
@@
The joint position bounds for the QP problem must be set in such a way that the robot joint limits are not exceeded. In our solution, we are inspired by the shaping policy to avoid joint limits in [8] to maximize reachability in the arm workspace. It consists of a flat region in most of the joint range. We replaced the original hyperbolic tangent functions with linear and constant functions to fulfill the constraint condition of the QP problem. Since the problem is solved at the velocity level, the joint position limits are expressed using the joint velocities for the joint $i$ as:
@@
IV Experiments and Results
We prepared seven experiments with a simulated iCub robot and three with a real one to validate the performance of the proposed solution and compare it with other solutions. An overview of the experiments can be seen in Tab. II.
@@ 13 Real Bimanual
IV-A Experiments in simulation
In simulation, we evaluated the performance of the controller (Sec. III-B) and the trajectory sampling (Sec. III-D). The first set of experiments aims to assess reachability, movement smoothness, and reference tracking, without obstacles nearby. Reachability is tested in a reaching experiment (Exp 1) with the targets taken from a grid of 3x3x3 positions in Cartesian space in random order and randomly switched two orientations. The positions are the Cartesian product of $x=$ [-0.23, -0.19, -0.15] m, $y=$ [0.11, 0.15, 0.19] m, $z=$ [0.08, 0.12, 0.16] m, and the orientations are $o_{1}=$ [-0.15, -0.79, 0.59, 3.06], $o_{2}=$ [-0.11, 0.99, 0.02, 3.14] in angle-axis notation. Smooth movement is evaluated in a point-to-point experiment (Exp 2), where the robot moves between two alternating poses $p_{1}=$ [-0.23, 0.26, 0.02, -0.15, -0.79, 0.59, 3.06] and $p_{2}=$ [-0.26, 0.03, 0.03, -0.11, 0.99, 0.01, 3.14], where the first three values are position coordinates in meters and the last four values are the orientation in angle-axis notation. Finally, a circular movement experiment (Exp 3) is chosen to validate the reference tracking. The overview of the experiments can be found in Tab. II.
In addition to the proposed solution, other alternatives and state-of-the-art methods are also evaluated. Specifically, we compare — 1) HARMONIOUS – the proposed solution in the single-arm version; 2) HARMONIOUS2 – the proposed solution in the dual-arm version when the goal of the other arm is to hold a specified position; 3) NEO – the controller by Haviland and Corke [2]; 4) React by Ngyuen et al. [3]; 5) Cart – Cartesian controller for the iCub robot [8], which is evaluated only in Exp 1 and Exp 2 as it primarily serves for point-to-point movements and does not feature avoidance constraints. Controllers 1–4 are evaluated with and without local trajectory sampling (LTS), described in Sec. III-D. Cart has its own local trajectory optimization.
@@ IV-A Experiments in simulation
TABLE III: Reachability experiment (Exp 1). Success rate of combined reachability in position and orientation. LTS stands for local trajectory sampling. Cart has its own local trajectory optimization. The success rate is computed from 5 repetitions (with a different order of the targets each time) of the experiment (in total 135 targets).
The results of the reachability experiment (Exp 1) are shown in Tab. III. In this experiment, our controller has the highest number of targets that are reached in 6D pose (i.e., position and orientation). The target is reached when the end effector is, at the same time, closer than 5 mm and 0.1 rad to the target. The success rate is computed from 5 repetitions (with a different order of the targets each time) of the experiment (in total 135 targets). The impact of the local trajectory sampling is visible as all controllers have better (or at least the same) reachability results with sampling than without sampling.
The results of the movement smoothness experiment (Exp 2) show the main impact of the local trajectory sampling (see Fig. 4(a)). Position sampling (Fig. 4(a) left) creates a bell-shaped Cartesian translation velocity profile typical of human arm motion [57]. The velocity profile of Cart is the most similar to the exact minimum-jerk movement, as it is specially optimized for it. Other versions are almost identical; only React does not reach zero velocity after 5 s. All versions without local sampling have snap onset and slow jerky decay. Orientation sampling (Fig. 4(a) right) creates a torque-minimal path for rotation as it travels along the straightest path of the rounded surface of a sphere and keeps the rotational velocity constant during movement [59], resulting in a visually smooth transition between two rotations. HARMONIOUS and NEO have a smooth angular velocity profile; React velocity profile is not smooth and does not reach zero after 5 s. Cart uses the same sampling for orientation as for position, and thus the resulting angular velocity profile is similar to that for translational velocity.
@@
The next experiment is a circular movement experiment with both arms in opposite directions (Exp 4). It shows how the dual-arm solution works, particularly arm prioritization and self-collision avoidance. As the reference controllers are for single arms only (or separately), only our solution is shown (without trajectory sampling). Figure 4(c) shows the trajectories of both end effectors making circular movements. In this case, the right arm was selected as the “primary arm”. Therefore, the right arm successfully tracks the circle all the time and the left arm prevents self-collision by keeping a safe distance instead of tracking the circle when the arms are too close.
@@
The last set of simulation experiments tests collision avoidance. For simplicity, the obstacle is virtual and is perceived without noise, omitting the influence of sensor data processing on the results. All three experiments are based on Exp 2 and differ only in the direction of the obstacle: 1) the obstacle moves along the movement, i.e., in the y-axis direction (Exp 5-1); 2) the obstacle moves towards the robot, i.e., in the x-axis direction (Exp 5-2); 3) the obstacle falls in the z-axis direction (Exp 5-3). In this case, Cart is not included in the comparison because it cannot handle obstacles. React is also omitted because it was outperformed already in the experiments without obstacles. More restrictive repulsive vectors were used instead of linear inequality constraints that allow motions in tangential directions. Therefore, we compare only HARMONIOUS and NEO. Figure 4(d) shows the fraction of targets reached during the experiment, the fraction of QP problems successfully solved during the experiments, and the distance between the obstacle and the closest point of the robot, where we assume that a distance of less than 25 mm is too dangerous. There is no significant difference between controllers when the obstacle moves along the robot’s movement (Exp 5-1). On the other hand, the obstacle that moves towards the robot (Exp 5-2) is problematic for NEO. Half of the time, the targets are not reached within a time limit and the controller cannot solve the QP problem in almost 40 % runs. Furthermore, small distances between the robot and the obstacle sometimes lead to collisions. Unlike NEO, HARMONIOUS has no problem with collisions or reaching in this experiment (see video S1 in Multimedia Materials). Our solution also works adequately in Exp 5-3. NEO did not reach all targets within the time limit, but the distance to an obstacle is safe.
IV-B Real robot experiments – collision avoidance
@@
The experiments with the real robot were prepared to evaluate the obstacle processing part of the solution, described in Sec. III-F. We show collision avoidance (visual and proximity modalities) and post-collision reaction (tactile modality) and their combination while the robot keeps position (Exp 11, see video S2 in Multimedia Materials) or moves back and forth between two points as in Exp 2 (Exp 12, see video S3 in Multimedia Materials). When visual obstacles are involved, the iCub gaze controller [60] is used to point the robot gaze to an appropriate position in the scene, thus compensating for the robot torso joint movements.
@@
Figure 5 graphically illustrates the experiment using snapshots. Figure 6 shows the operation of the main components of HARMONIOUS on plots, focusing on the left arm of the robot (the “primary arm” in these experiments). Note that obstacles at or near the robot hand (visual obstacles near the hand or forearm, proximity at the back of the hand, or touch at the hand or forearm) inevitably compromise the task—staying close to the target—whereas visual obstacles or touch at more proximal body parts give the robot the possibility to exploit its kinematic redundancy and fulfill the task while simultaneously avoiding contacts. Parts of Exp 11—robot tasked with keeping the end effector position while complying with constraints from obstacles spawned by the presence of the human—are shown in Fig. 6(a). In Fig. 6(a) (top), the robot starts in a steady state without obstacles around characterized by a zero target distance (first two seconds, green line). Once a proximity obstacle is detected, the end effector is moving away until the proximity obstacle constraint disappears from the controller and the target position is reached again. Physical contact of the human with the robot hand (around 5 s) and the forearm (around 7 and 9 s) generate constraints and trigger avoidance. Tactile collision with upper arm (12-13 s) and torso (14-15 s) generate avoidance but do not significantly compromise the task, as the robot kinematic redundancy is exploited by the controller.
@@
In Fig. 6(a) (center), the first 10 seconds are very similar to the top one. Touch on the forearm causes deviation from the target while contact with the upper arm is evaded while simultaneously staying close to the target. Touch at the hand (5 s) and proximity signal at the hand (7 s) cause deviation from the target. At around 12 s, a proximity obstacle at the hand is evaded, but then avoidance stops as the robot is blocked by contact with the forearm. Once the touch disappears, the proximity collision avoidance continues. Finally, when the proximity obstacle disappears, a tactile stimulus on the hand causes a fast movement towards the target position.
@@
IV-C Real robot – bimanual task
We prepared a bimanual task experiment (Exp 13, see video S4 in Multimedia Materials) in which the robot holds a sponge in its hands (see Fig. 7(a)). The target is prescribed for the “primary arm” (left) and an additional constraint is introduced to keep the sponge between the two hands. At the same time, dynamic obstacles spawn additional constraints and avoidance. For this experiment, tactile inputs from the robot hands (palms) are ignored (the robot holds the sponge). This experiment aims to verify that the dual-arm version works and that other task constraints, such as the relative distance vector between the end effectors in this case, can be easily incorporated into the problem definition (the relative position constraint described in Sec. III-J).